# SOP-Bench: Complex Industrial SOPs for Evaluating LLM Agents

### v1

@@ 3.1.1 Dataset Schema Generation
    • smoking_statusstring, multiple choice for current smoking habits, Choices = Never, Former, Current
  • SOP Generated Without data schema entry in context
@@
    • ... Calculate the smoking risk index using preferred methodology ...
  • SOP Generated With data schema entry in context
@@
    • ... Calculate the smoking risk index (Never, Former, Current) ...
3.1.2 SOP Document Generation
The Standard Operating Procedure (SOP) document captures all operational steps required to complete a task. Using <Business Task, Task Context, Dataset Schema> (Appendix C.1), we generate a structured SOP that outlines procedure objectives, prerequisites, input requirements, detailed instructions, and decision logic in matural language. The inclusion of task context and a human-verified dataset schema forms a unified semantic scaffold that grounds the SOP in domain-specific structure and preserves causal and operational dependencies. This ensures the resulting SOPs are clear, consistent, and aligned with real-world workflows. Human oversight is used to validate domain standards, task logic, and correct inconsistencies or hallucinations, ensuring the SOPs are both practically executable and contextually appropriate for industrial use.
3.1.3 Dataset Generation
@@
We then utilize <Business Task, Task Context, Dataset Schema, SOP>(Appendix  C.3) to generate the underlying dataset that serves as the foundation for SOP execution and evaluation. The dataset represents full procedural context as well as structured inputs and outputs for every tool/API invocation, task parameters, decision pathways and final outcomes. By generating the dataset using dataset schema and SOP, we ensure that the dataset is contextually grounded. Human oversight is used to ensure each datapoint’s logic is unambiguous, the dataset contains appropriate data-types and accurately reflects the SOP logic. The dataset is constructed to include datapoints that challenge the reasoning, problem-solving, compliance, error-handling and tool selection capabilities of LLM agents by including both positive and negative decision pathways as well as edge conditions and task failure scenarios.
3.1.4 API & Tool Specification Generation
@@
To enable LLM agents to dynamically retrieve task-relevant information, we use the generated datasets to define APIs that map to the operational flow of the SOP. Using <Task Context, Dataset, SOP> (Appendix C.4), we generate APIs specifying required inputs (e.g., patient_name) and expected outputs (e.g., patient_id) (Figure  4 in Appendix B). To inform the agent of available tools at each step, we generate corresponding tool specifications using <Task Context, Dataset, SOP, APIs>, also shown in Figure 4 in Appendix B). Human experts validate that APIs and ToolSpecs align with the dataset schema and access only relevant columns, ensuring semantic consistency and operational validity. Here is an example which demonstrates that without clear specifications, the LLM lacks the necessary context to infer input parameters and expected outputs.
  • API Specification
    • getPatientID – retrieves a unique patient identifier based on a provided name  4
  • Tools Code Generation Without API in context
    • def getPatientID(self):
    • return {}
  • Tools Code Generation With API in context
    • def getPatientID(self, patient_name):
    • # Calls internal hospital system to fetch patient ID
@@
    • return self.patientDB.lookupIDByName(patient_name)
In contrast, explicit API specs provide structure and semantic clarity, enabling the LLM to generate more accurate, robust, and callable tool implementations that integrate correctly into the larger system.
3.1.5 Tools Code Generation
To generate executable code for LLM agents to process tasks, we used <Dataset, APIs> (Appendix C.5) to produce tool code reflecting the logic needed for each procedural step. The dataset snapshot and API specs provide input-output contracts as well as data characteristics, ensuring generated code aligns with operational context and system requirements. By processing all APIs together, the LLM identifies reusable patterns such as validations, conditionals, error handling and distills them into consistent templates to improve code efficiency and reliability.
3.1.6 ** Adding complexity in generated components **
In our benchmark generation workflow, SOP-level complexity is primarily introduced through branching logic (e.g., “If a condition fails, choose path P; else path Q”), requiring AI agents to reason over multiple potential execution paths. Each SOP also consists of a set of tools that the agents must use to execute certain functionalities such as fetching data, verifying status, etc.
To evaluate whether agents can address real-world ambiguities while maintaining high performance when faced with confusing and overlapping instructions in a SOP, we applied a strategy of rewriting a SOP to include additional redundant details that are typically related to the business task, but not directly relevant to executing the core steps of the SOP. These additions mimic real-world documentation noise, partially obfuscating the key steps in the SOP with contextual background, tangential best practices and outdated alternatives, which agents must learn to filter out. In parallel, with some SOPs, we introduced semantically similar but functionally redundant tools, creating scenarios where the agents must critically evaluate each option rather than relying on surface cues.
Together, these interventions, induced either by LLMs via prompting or manual authoring by humans, and validated through human oversight, allow agent builders to rigorously assess their agent’s tools selection behavior, identify failure modes, and iteratively refine both the tool definitions and the agent’s decision-making policy for truly robust SOP comprehension and tool selection.
@@ 3.1.5 Tools Code Generation
Table 3: Characteristics of Benchmark SOPs in SOP-Bench. $n$ = number of tasks per SOP; $t$ = number of tools available; $\tau$ = token count, using GPT 3.5 turbo [14]; $\mathbb{C_{H}}$ = human-perceived complexity, average of three domain experts of three dimensions, ease of understanding, ambiguity or implicit domain knowledge and reasoning complexity, on a scale of 1 - 10; $\mathbb{C}_{llm}$ = complexity estimated by Claude 3.5 (scale 1-10, calibrated against Berkeley function calling tasks as baseline 2)
@@
Average 181 11 132
Total 1,811 107
3.2 SOP Descriptions
@@
SOP-Bench consists of 10 diverse generic SOPs instantiated in 1800+ tasks that were designed to represent realistic industrial procedures. Each SOP presents unique challenges that test different aspects of agent capabilities, from complex decision-making to tool usage and structured output generation. These SOPs expose agents to a broad spectrum of business modalities that include classification such as categorizing email intent, structured verification based modalities such as business verification, industrial processes such as aircraft pre-flight inspection as well as generic business assignments such as annotation, content flagging, customer service, etc. Each of these SOPs help agent builders to explore different forms of reasoning, compliance-handling, nuanced intent interpretation and procedural task planning and execution. See appendix B for details. The detailed stats are listed in table 3, where we enumerate the topics, domain, task count ($n$), API count ($t$), SOP token count ($\tau$) and rate of complexity levels by we ask three human domain experts to rate the SOPs on (1) ease of understanding (2) ambiguity or implicit domain knowledge required to comprehend the SOP, and (3) reasoning complexity. We average the scores and present in the table. We also ask Claude 3.5 v2 to rate the SOPs on complexity on a scale of 1 - 10.
4 Agents for automating SOPs
@@
To evaluate the effectiveness of LLM-based agents in executing synthetically generated SOPs that mimic their real-world counterparts, we implement two complementary agent architectures in SOP-Bench: (a) Function Calling (FC) Agent and (b) ReAct Agent. Both agents utilize Anthropic’s Claude 3.5 Sonnet v2 model, leveraging Claude’s native tool-calling interface to dynamically invoke external functions/APIs during task execution. More details on these two agents are in Appendix  A.
@@
We evaluate both agents on SOP-Bench using standardized scripts that: (1) load SOP documents and structured task datasets; (2) convert each task into natural language instructions; and (3) execute agents while logging outputs, tool traces, and reasoning steps. To simulate API calls, each dataset includes precomputed inputs and outputs for every tool call, stored as columns. These mocks replace live APIs to enable stable, reproducible evaluation without runtime variability. In deployment, they would be swapped for real APIs. For some SOPs, we also compare intermediate decisions, based on API responses, with expected ground truth to assess multi-step reasoning. Though not returned by tools, these values are logged for evaluating agent reasoning. This setup supports controlled comparisons of agent behavior, tool use, and LLM performance across diverse SOPs.
5 Evaluating Agents with SOP-Bench

### v2

@@ 3.1.1. Dataset Schema Generation (LLM-Assisted, Human-Validated)
  • Schema Entry: smoking_statusstring, Choices = Never, Former, Current
@@
  • Without Schema : Calculate smoking risk using preferred methodology
@@
  • With Schema : Calculate smoking risk (Never, Former, Current)
3.1.2. SOP Refinement (Human-Authored, AI-Refined)
Our workflow next uses the first version of the SOP, along with <Business Task, Task Context, Dataset Schema> (Appendix C.1) to refine the SOP and make sure it is consistent with the data schema. The resulting SOP outlines procedure objectives, prerequisites, input requirements, detailed instructions, and decision logic in natural language. Human oversight validates domain standards, task logic, and corrects inconsistencies, ensuring SOPs are executable and contextually appropriate for industrial use.
3.1.3. Dataset Generation (AI-Generated, Human-Validated)
@@
Using <Business Task, Task Context, Dataset Schema, SOP> (Appendix C.3), AI generates datasets representing full procedural context: structured inputs/outputs for every tool invocation, task parameters, decision pathways, and final outcomes. Human experts validate that each datapoint’s logic is unambiguous, data types are appropriate, and the dataset accurately reflects SOP logic. The dataset includes challenging scenarios: positive/negative decision pathways, edge conditions, and task failure scenarios to test agent reasoning, compliance, error-handling, and tool selection.
3.1.4. API & Tool Specs Generation (AI-Generated, Human- Validated)
@@
Using <Task Context, Dataset, SOP> (Appendix C.4), our workflow generates APIs specifying required inputs (e.g., patient_name) and expected outputs (e.g., patient_id), along with corresponding tool specifications (Figure 4). Human experts validate that APIs and ToolSpecs align with the dataset schema and access only relevant columns. Example:
  • API Spec: getPatientID – retrieves patient identifier from name
  • Without API : def getPatientID(self): return {}
@@
  • With API : def getPatientID(self, patient_name): return self.patientDB.lookupIDByName(patient_name)
3.1.5. Tools Code Generation (AI-Generated, Human-Validated)
Using <Dataset, APIs> (Appendix C.5), Our workflow next generates executable tool code reflecting procedural logic. The dataset and API specs provide input-output contracts and data characteristics, ensuring code aligns with operational requirements. Human experts validate code correctness and consistency by executing them.
@@ 3.1.5. Tools Code Generation (AI-Generated, Human-Validated)
Table 3. Characteristics of Benchmark SOPs in SOP-Bench. $n$ = number of tasks per SOP; $t$ = number of tools available; $\tau$ = token count, $\mathbb{C_{H}}$ = human-perceived complexity, average of three domain experts of three dimensions, ease of understanding, ambiguity or implicit domain knowledge and reasoning complexity, on a scale of 1 - 10; $\mathbb{C}_{llm}$ = complexity estimated by Claude 3.5 (scale 1-10, calibrated against Berkeley function calling tasks as baseline 2)
@@
Referral Abuse v1 Trust & Safety 200 3 1502 6.33 7
Referral Abuse v2 Trust & Safety 200 6 2576 8.67 9
Traffic Spoofing Fraud Detection 200 6 854 6.83 8
Average 185 9 1394 6.96 8.31
Total 2411 122 18120
3.2. Human Authoring and Validation
Domain experts perform five critical validation steps: (1) SOP Authoring: experts draft initial SOPs based on real industrial procedures, providing business context and task requirements; (2) Schema Validation: experts verify that AI-generated schemas capture essential task components, implicit domain knowledge, and compliance constraints; (3) SOP Refinement: experts review AI-refined SOPs for domain accuracy, logical consistency, and alignment with industrial standards; (4) Dataset Validation: experts validate each generated datapoint for logical correctness, appropriate data types, and coverage of edge cases; (5) API & Code Validation: experts verify that generated APIs and executable code align with dataset schemas and operational requirements.
To quantify benchmark difficulty, three domain experts independently rated each SOP across three dimensions: ease of understanding, ambiguity/implicit knowledge requirements, and reasoning complexity (scale 1-10). These ratings, averaged to produce $\mathbb{C}_{H}$ in Table 3, reflect human-perceived complexity.
3.3. SOP Descriptions
@@
SOP-Bench consists of 12 diverse SOPs instantiated in 2,411 tasks representing realistic industrial procedures. Each SOP tests different agent capabilities: complex decision-making, tool usage, and structured output generation across business modalities including classification (email intent), structured verification (business verification), industrial processes (aircraft inspection), and generic assignments (annotation, content flagging, customer service). Table 3 shows task count ($n$), API count ($t$), SOP token count ($\tau$), and complexity ratings by three human domain experts ($\mathbb{C}_{H}$) and Claude 3.5 v2 ($\mathbb{C}_{llm}$).
3.3.1. Controlled Complexity Variation
To evaluate how procedural complexity affects agent performance, we include two versions of the referral abuse detection SOP. Version 1 (v1) implements a foundational fraud detection workflow with 3 tools and a simple three-outcome decision model based on current account indicators. Version 2 (v2) introduces significantly greater complexity through temporal pattern analysis, historical violation context, and risk-based enforcement, requiring 6 tools and producing 7 graduated enforcement actions. This controlled complexity variation enables systematic evaluation of how agents handle increasingly sophisticated reasoning requirements.
3.3.2. Real-World Complexity in SOP-Bench
To mirror authentic industrial conditions, human experts and AI collaboratively introduce controlled complexity at multiple levels across different SOPs. At the SOP-level, procedures contain branching logic, redundant details, tangential information, and ambiguous instructions that agents must filter—reflecting the reality that real-world SOPs are rarely perfectly structured. At the tool-level, we include semantically similar but functionally redundant tools to challenge agents’ critical evaluation and tool selection capabilities. At the data-level, we introduce distractor variables that appear relevant but are not required for task completion.
Specific examples include: (1) Video Annotation and Video Classification contain 26 and 25 tools respectively, with multiple distractor tools that perform similar functions (e.g., different object detection methods, overlapping segmentation approaches), requiring agents to identify the correct tool sequence from numerous plausible alternatives; (2) Patient Intake includes distractor input variables such as patient demographic details and medical history fields that are provided but not necessary for certain decision pathways, testing whether agents can distinguish essential from supplementary information; (3) Content Flagging and Customer Service embed ambiguous conditional logic where threshold values and decision criteria require careful interpretation of SOP text rather than explicit enumeration. These interventions, validated through human oversight, rigorously assess tool selection behavior, information filtering, and reasoning under ambiguity. See Appendix B for detailed descriptions of all SOPs.
4. Agents for automating SOPs
Table 4. Claude 4 Opus: Function Calling vs ReAct Agent Performance, ranked in descending order of ReAct TSR.
SOP Topic Function Calling ReAct
ECR C-TSR TSR Time (s) ECR C-TSR TSR Time (s)
Patient Intake 1.00 0.92 0.92 70.33 1.00 1.00 1.00 102.28
Email Intent 1.00 0.98 0.98 25.24 0.99 0.99 0.98 31.26
Referral Abuse Detection Easy 1.00 0.95 0.95 53.99 1.00 0.97 0.97 66.45
Aircraft Inspection 1.00 0.45 0.45 83.11 1.00 0.97 0.97 87.46
Referral Abuse Detection Hard 1.00 0.98 0.98 81.79 0.99 0.96 0.96 89.16
Video Classification 1.00 0.79 0.79 78.69 1.00 0.94 0.94 62.89
Dangerous Goods 1.00 0.73 0.73 41.15 1.00 0.83 0.83 28.02
Warehouse Package Inspection 1.00 0.56 0.56 47.50 1.00 0.65 0.65 76.02
Traffic Spoofing Detection 1.00 0.79 0.79 25.23 0.99 0.61 0.60 130.20
Customer Service 1.00 0.56 0.56 72.83 1.00 0.56 0.56 145.43
Video Annotation 1.00 0.37 0.37 71.04 1.00 0.38 0.38 73.97
Know Your Business 0.99 0.44 0.43 159.84 1.00 0.31 0.31 110.50
Content Flagging 1.00 0.36 0.36 49.93 0.99 0.28 0.27 138.08
Average 1.00 0.68 0.68 66.21 1.00 0.73 0.72 87.82
Standard Error $\epsilon$ $\pm$0.07 $\pm$0.07 $\pm$9.57 $\pm$0.08 $\pm$0.08 $\pm$10.30
@@ 4. Agents for automating SOPs
To evaluate the effectiveness of LLM-based agents in executing industry grade SOPs, we implement two complementary agent architectures in SOP-Bench: (a) Function Calling (FC) Agent and (b) ReAct Agent. These implementations serve as baseline agents to demonstrate SOP-Bench’s evaluation capabilities and are not intended to represent optimal architectures. We encourage the research community to develop and publish more sophisticated agent designs using SOP-Bench as the evaluation framework. Both baseline agents utilize LLM’s native tool-calling interface to dynamically invoke tools/APIs during task execution. More details on these two agents are in Appendix  A.
FC Agent is a prompt-driven, lightweight execution engine leveraging native tool calling capabilities of LLMs. It accepts SOP text and task input, maintains a conversation loop with the LLM, and processes tool calls iteratively (up to 10 iterations). When the model triggers a tool use event, the agent executes the corresponding function via a ToolManager and returns results to the model. Final outputs are enclosed in XML tags for structured parsing.
ReAct Agent is built on a custom ReAct framework implementation (Yao et al., 2023), designed for SOPs with complex branching logic and multi-step reasoning. It interleaves reasoning traces (Thoughts) with tool calls (Actions), feeding Observations back into the reasoning loop for up to 15 iterations. The agent enforces that at least one tool is executed before conclusions, maintaining comprehensive execution traces including intermediate steps and tool call records. More details on these two agents are in Appendix  A.
@@
We evaluate both agents on SOP-Bench using standardized scripts that: (1) load SOP documents and structured task datasets; (2) convert each task into natural language instructions; and (3) execute agents while logging outputs, tool traces, and reasoning steps. To simulate API calls, each dataset includes precomputed inputs and outputs for every tool call, stored as columns. These mocks replace live APIs to enable stable, reproducible evaluation without runtime variability. In deployment, they would be swapped for real APIs.
4.1. Evaluation Methodology